The half-life of a radioactive isotope is 2.5 min.
Answer #1

Number of atoms, Nt = Noe-(ln2)t/t1/2

So, N15 = (48 * 1032) * e-(ln2 * 15 / 2.5) = 7.5 * 1031 atoms

Decay constant, λ = ln2 / t1/2 = ln2 / 2.5 = 0.28 min-1

Initial activity, Ao = λNo = 0.28 * (7.5 * 1031) = 2.1 * 1031 decays/min